# Multi-Robot Object SLAM using Distributed Variational Inference

### v1

@@ Abstract
Multi-robot simultaneous localization and mapping (SLAM) enables a robot team to achieve coordinated tasks relying on a common map. However, centralized processing of robot observations is undesirable because it creates a single point of failure and requires pre-existing infrastructure and significant multi-hop communication throughput. This paper formulates multi-robot object SLAM as a variational inference problem over a communication graph. We impose a consensus constraint on the objects maintained by different nodes to ensure agreement on a common map. To solve the problem, we develop a distributed mirror descent algorithm with a regularization term enforcing consensus. Using Gaussian distributions in the algorithm, we derive a distributed multi-state constraint Kalman filter (MSCKF) for multi-robot object SLAM. Experiments on real and simulated data show that our method improves the trajectory and object estimates, compared to individual-robot SLAM, while achieving better scaling to large robot teams, compared to centralized multi-robot SLAM. Code is available at https://github.com/intrepidChw/distributed_msckf.
@@ I INTRODUCTION
This paper considers a multi-robot landmark-based SLAM problem. We develop an approach for distributed Bayesian inference over a graph by formulating a mirror descent algorithm [7] in the space of probability density functions and introducing a regularization term that couples the estimates of neighboring nodes. Our formulation allows joint optimization of common variables (e.g., common landmarks among the robots) and local optimization of others (e.g., private robot trajectories). As a result, each node is able to keep a distribution only over its variables of interest, enabling both efficient storage and communication. By using Gaussian distributions in the mirror descent algorithm, we derive a distributed version of the widely used multi-state constraint Kalman filter (MSCKF) [8] with an additional averaging step to enforce consensus for the common variables. We apply the distributed MSCKF algorithm to collaborative object SLAM using only stereo cameras at each robot. In the prediction step, each robot estimates its trajectory locally using visual odometry. In the update step, the robots correct their trajectory estimates using both visual features and object detections. As is common in the MSCKF, we avoid keeping visual landmarks in the robot states using a null-space projection step. However, object landmarks are kept as a map representation at each robot and are shared among the robots during the consensus averaging step to collaboratively estimate consistent object maps. In short, each robot estimates its pose trajectory and an object map locally but communicates with its neighbors to reach agreement on the object map across the robots.
The contributions of this work are summarized as follows.
@@
  • We demonstrate multi-robot object SLAM using stereo camera measurements for odometry and object detection on both real and simulated data to show that our method improves the overall accuracy of the trajectories and object maps of robot teams, compared to individual-robot SLAM, while achieving better scaling to large robot teams, compared to centralized multi-robot SLAM.
@@ II RELATED WORK
II-A Multi-robot pose graph optimization
Tian et al. [15] propose a certifiably correct pose graph optimization (PGO) method with a novel Riemannian block coordinate descent (RBCD) that can operate in a distributed setting. Considering both localization and mapping, Cunningham et al. [16, 17] extend Smoothing and Mapping (SAM) [18] by introducing a Constrained Factor Graph that enforces consistent estimations of common landmarks among robots. Choudhary et al. [19] developed a two-stage approach using Successive Over-Relaxation and Jacobi Over-Relaxation to split the computation among the robots. MR-iSAM2 [20] extends incremental smoothing and mapping (iSAM2) [10] by introducing a novel data structure called mult-root Bayes tree. Tian et al. [21] investigate the relations between Hessians of Riemannian optimization and Laplacians of weighted graphs and propose a communication-efficient multi-robot optimization algorithm with a central server performing approximate second-order optimization.
@@ II-B Multi-robot filtering
Roumeliotis and Bekey [28] showed that the Kalman filter equations can be written in decentralized form, allowing decomposition into smaller communicating filters at each robot. Thrun et. al [29] presented a sparse extended information filter for multi-robot SLAM, which actively removes information to ensure sparseness at the cost of approximation. With nonlinear motion and observation models, a decentralized extended Kalman filter (EKF) has an observable subspace of higher dimension than the actual nonlinear system and generates unjustified covariance reduction [30]. Huang et al. introduced observability constraints in EKF [30] and unscented smoothing [31] algorithms to ensure consistent estimation. Gao et al. [32] use random finite sets to represent landmarks at each robot and maintain a probability hypothesis density (PHD). The authors prove that geometric averaging of the robot PHDs over one-hop neighbors leads to convergence of the PHDs to a global Kullback-Leibler average, ensuring consistent maps across the robots. Zhu et al. [33] propose a distributed visual-inertial cooperative localization algorithm by leveraging covariance intersection to compensate for unknown correlations among the robots and deal with loop-closure constraints.
Compared to existing work, our contribution is to derive a fully distributed filtering-based system for landmark-based SLAM from a constrained variational inference perspective. Our formulation makes a novel connection to distributed mirror descent and enables robots to achieve landmark consensus efficiently with one-hop communication only and without sharing private trajectory information.
@@ III PROBLEM STATEMENT
The robots communicate over a network represented as a connected undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with nodes $\mathcal{V}=\{1,\ldots,n\}$ corresponding to the robots and edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ specifying robot pairs that can exchange information. For example, $(i,j)\in\mathcal{E}$ indicates that robot $i$ and $j$ can exchange information. Let $A\in\mathbb{R}^{n\times n}$ be a doubly stochastic weighted adjacency matrix of $\mathcal{G}$ such that $A_{ij}>0$ if $(i,j)\in\mathcal{E}$ and $A_{ij}=0$ otherwise. Also, let $\mathcal{N}_{i}:=\{j\in\mathcal{V}|(j,i)\in\mathcal{E}\}\cup\{i\}$ denote the set of one-hop neighbors of robot $i$. We consider the following problem.
@@
Given control inputs $\mathbf{u}_{i}=[\mathbf{u}_{i,0}^{\top},\ldots,\mathbf{u}_{i,T-1}^{\top}]^{\top}$ and observations $\mathbf{z}_{i}:=[\mathbf{z}_{i,0}^{\top},\ldots,\mathbf{z}_{i,T}^{\top}]^{\top}$ at each robot $i$, estimate the robot states $\mathbf{s}_{i}:=[\mathbf{s}_{i,0}^{\top},\ldots,\mathbf{s}_{i,T}^{\top}]^{\top}$ collaboratively by communicating only with one-hop neighbors according to the structure of the graph $\mathcal{G}$.
@@ IV-B Distributed mirror descent
The variational inference problem in (5) can be solved using the mirror descent algorithm [7]. Mirror descent is a generalization of projected gradient descent that performs projection using a generalized distance (Bregman divergence), instead of the usual Euclidean distance, to respect the geometry of the constraint set $\mathcal{Q}_{i}$. Since $\mathcal{Q}_{i}$ is a space of PDFs, a suitable choice of Bregman divergence is the KL divergence. Starting with a prior PDF $q_{i}^{(0)}(\mathbf{s}_{i})$, the mirror descent algorithm performs the following iterations:
@@
where $\delta c_{i}/\delta q_{i}(q_{i}^{(k)})$ is the Fréchet derivative of $c_{i}(q_{i})$ with respect to $q_{i}$ evaluated at $q_{i}^{(k)}$ and $\alpha_{k}>0$ is a step size.
Note that the optimizations (6) at each robot $i$ are completely decoupled and, hence, each robot would be estimating its own density over the common landmarks $\mathbf{y}$. To make the estimation process collaborative, the regularization term $\operatorname*{KL}(q_{i}||q_{i}^{(k)})$ in (6) should require that the PDF $q_{i}$ of robot $i$ is also similar to the priors $q_{j}^{(k)}$ of its neighbors $\mathcal{N}_{i}$ rather than its own prior $q_{i}^{(k)}$ alone. This idea was introduced in the distributed mirror descent algorithm in [35], where a weighted sum of KL divergences with respect to each neighbor is used for regularization. In our case, the PDFs $q_{j}^{(k)}(\mathbf{s}_{j})=q_{j}^{(k)}(\mathbf{x}_{j},\mathbf{y})$ are not defined over the same set of variables since each robot $j$ is estimating its own private state $\mathbf{x}_{j}$ as well. To enforce consensus only on the common state $\mathbf{y}$, the KL divergence term in (6) can be decomposed as a sum of marginal and conditional terms:
@@ IV-C Linear Gaussian case
Denoting the density and distribution at iteration $k$ as $q_{i}^{(k)}(\mathbf{s}_{i})$ and $\mathcal{N}(\boldsymbol{\mu}_{i,(k)},\Sigma_{i,(k)})$, the distributed variational inference update in (9) is computed in the following propositions.
@@
As shown in [36, Chapter 3.3], considering only two consecutive time steps in the lifted form in (11) leads to a Kalman filter. Using the result in Proposition 3, we obtain a distributed Kalman filter that incorporates the consensus averaging step in (15).
@@ V-B Consensus averaging
which is the same as (15) except that we only consider one time step $t$ here. Then, we need to reconstruct the new joint distribution $\mathcal{N}(\bar{\boldsymbol{\mu}}_{i,t-1},\bar{\Sigma}_{i,t-1})$. Since we store the historical camera poses as $SE(3)$ matrices, Proposition 2 can not be applied directly. Following [36, Chapter 7.3.1], we define a Gaussian distribution over a historical camera pose $\underline{T}_{i,k},\ k=t-c,\cdots,t-1$ by adding a perturbation $\epsilon_{i,k}$:
@@
where $(\cdot)^{\wedge}$ defined in [36, Chapter 7.1.2] converts from $\mathbb{R}^{6}$ to a $\mathbb{R}^{4\times 4}$ twist matrix. The estimated covariance $\Sigma_{i,t-1}$ already takes account of both the poses and landmarks. For consistency of notation with Proposition 2, we denote the mean of the pose perturbation as $\boldsymbol{\mu}^{\mathbf{x}}_{i,t-1}=\mathbf{0}_{6c}$. After averaging and reconstructing the new joint distribution, $\bar{\boldsymbol{\mu}}^{\mathbf{x}}_{i,t-1}$ may be non-zero, so we need to correct the camera poses as follows:
@@ V-C State propagation
We derive a general odometry propagation step for the MSCKF algorithm, thus not necessarily requiring IMU measurements and enabling vision-only propagation. We assume an odometry algorithm (e.g., libviso2 [37]) provides a relative pose measurement $\delta T_{i,t-1}$ between the previous frame at time $t-1$ and the current frame at time $t$. The state at each robot $i$ is propagated as:
@@
where $Ad(\cdot)$ is the adjoint of an $SE(3)$ matrix [36, Chapter 7.1.4], $\mathbf{e}_{6n}\in\mathbb{R}^{6n+3m_{t}}$ is a vector with the $6n$-th element as $1$ and the rest as 00, and $W_{i}\in\mathbb{R}^{6\times 6}$ is the odometry measurement covariance.
@@ V-D State update
The MSCKF update step follows the previous work [38]. We first define several residuals. The camera pose residual is a perturbation $\underline{\epsilon}_{i,k}$ that transforms the estimated pose $T_{i,k}$ to the true pose $\underline{T}_{i,k}$, i.e. $\underline{T}_{i,k}=T_{i,k}\exp(\underline{\epsilon}_{i,k}^{\wedge})$. The landmark residual is the difference between true position $\underline{\mathbf{p}}$ and the estimated position $\mathbf{p}$, i.e., $\tilde{\mathbf{p}}=\underline{\mathbf{p}}-\mathbf{p}$, where $\mathbf{p}$ is the position mean if the landmark is in the state or the result triangulation if it is not. When robot $i$ receives the $k$-th geometric feature observation at time $t$, denoted as $\mathbf{z}^{g}_{i,t,k}$, we linearize the observation model around the current error state $\tilde{\mathbf{s}}_{i,t,k}$ (composed of both pose and landmark residuals) and the feature position residual $\tilde{\mathbf{p}}^{g}_{i,k}$:
@@
where $\hat{\mathbf{z}}^{g}_{i,t,k}$ is the predicted observation based on the observation model and $H^{\mathbf{s},g}_{i,t,k}$ and $H^{\mathbf{p},g}_{i,t,k}$ are Jacobians. Then, we left-multiply by the nullspace $N_{i,t,k}$ of $H^{\mathbf{p},g}_{i,t,k}$ to remove the effect of $\tilde{\mathbf{p}}^{g}_{i,k}$:
@@
Concatenating $\mathbf{r}^{g,0}_{i,t,k}$, $N_{i,t,k}^{\top}H^{\mathbf{s},g}_{i,t,k}$ for all $k$, we get an overall geometric feature residual $\mathbf{r}^{g}_{i,t}$ and Jacobian $H^{g}_{i,t}$.
Similarly, we linearize the observation model for the object features $\mathbf{z}^{o}_{i,t,k}$:
@@
and concatenate $\mathbf{r}^{o}_{i,t,k}$ and $H^{\mathbf{s},o}_{i,t,k}$ for all $k$ to get an overall object observation residual $\mathbf{r}^{o}_{i,t}$ and Jacobian $H^{o}_{i,t}$.
Finally, by concatenating the residuals and Jacobians of both geometric and object features, we get the overall residual $\mathbf{r}_{i,t}=[\mathbf{r}^{g\top}_{i,t}\ \mathbf{r}^{o\top}_{i,t}]^{\top}$ and Jacobian $H_{i,t}=[H^{g\top}_{i,t}\ H^{o\top}_{i,t}]^{\top}$ and compute the Kalman gain:
$K_{i,t}=\Sigma^{+}_{i,t}H^{\top}_{i,t}(H_{i,t}\Sigma^{+}_{i,t}H^{\top}_{i,t}+I\otimes V_{i})^{-1},$ (25)
where $\Sigma^{+}_{i,t}$ comes from the propagation step and $K_{i,t}$ can be split to $K^{\mathbf{x}}_{i,t}$ and $K^{\mathbf{y}}_{i,t}$ related to $\mathbf{x}_{i,t}$ and $\mathbf{y}_{i,t}$ respectively. The landmark mean, camera poses, and the entire covariance are updated as:

### v2

@@ Abstract
Multi-robot simultaneous localization and mapping (SLAM) enables a robot team to achieve coordinated tasks by relying on a common map of the environment. Constructing a map by centralized processing of the robot observations is undesirable because it creates a single point of failure and requires pre-existing infrastructure and significant communication throughput. This paper formulates multi-robot object SLAM as a variational inference problem over a communication graph subject to consensus constraints on the object estimates maintained by different robots. To solve the problem, we develop a distributed mirror descent algorithm with regularization enforcing consensus among the communicating robots. Using Gaussian distributions in the algorithm, we also derive a distributed multi-state constraint Kalman filter (MSCKF) for multi-robot object SLAM. Experiments on real and simulated data show that our method improves the trajectory and object estimates, compared to individual-robot SLAM, while achieving better scaling to large robot teams, compared to centralized multi-robot SLAM.
@@ I INTRODUCTION
This paper considers a multi-robot landmark-based SLAM problem. We develop an approach for distributed Bayesian inference over a graph by formulating a mirror descent algorithm [7] in the space of probability density functions and introducing a regularization term that couples the estimates of neighboring nodes. Our formulation allows joint optimization of common variables (e.g., common landmarks among the robots) and local optimization of others (e.g., private robot trajectories). As a result, each node keeps a distribution only over its variables of interest, enabling both efficient storage and communication. By using Gaussian distributions in the mirror descent algorithm, we derive a distributed version of the widely used multi-state constraint Kalman filter (MSCKF) [8] with an additional averaging step to enforce consensus for the common variables. We apply our distributed MSCKF algorithm to collaborative object SLAM using only stereo camera observations at each robot. In the prediction step, each robot estimates its trajectory locally using visual odometry. In the update step, the robots correct their trajectory estimates using both visual features and object detections. As common in the MSCKF, we avoid keeping visual landmarks in the state using a null-space projection step. However, object landmarks are kept as a map representation for each robot and are shared among the robots during the consensus averaging step to collaboratively estimate consistent object maps. In short, each robot estimates its trajectory and an object map locally but communicates with its neighbors to reach agreement on the object maps across the robots. Our contributions are summarized as follows.
@@
  • We demonstrate multi-robot object SLAM using stereo camera measurements for odometry and object detection on both real and simulated data to show that our method improves the overall accuracy of the trajectories and object maps of robot teams, compared to individual-robot SLAM, while achieving better scaling to large robot teams, compared to centralized multi-robot SLAM [^1] Code is available at https://github.com/intrepidChw/distributed_msckf..
@@ II RELATED WORK
II-A Multi-robot bluefactor graph optimization
Factor graph methods formulate SLAM as an optimization problem over a bipartite graph of variables to be estimated and factors relating variables and measurements via error functions. Tian et al. [15] propose a certifiably correct pose graph optimization (PGO) method with a novel Riemannian block coordinate descent (RBCD) that operates in a distributed setting. Cunningham et al. [16, 17] extend smoothing and mapping (SAM) [18] by introducing a constrained factor graph that enforces consistent estimates of common landmarks among robots. Choudhary et al. [19] developed a two-stage approach using successive over-relaxation and Jacobi over-relaxation to split the computation among the robots. MR-iSAM2 [20] extends incremental smoothing and mapping (iSAM2) [10] by introducing a novel data structure called mult-root Bayes tree. Tian et al. [21] investigate the relation between Hessians of Riemannian optimization and Laplacians of weighted graphs and design a communication-efficient multi-robot optimization algorithm performing approximate second-order optimization.
@@ II-B Multi-robot filtering
Filtering methods, such as the Kalman filter, offer a computationally lightweight alternative to factor graph optimization by performing incremental prediction and update steps that avoid a large number of iterations. Roumeliotis and Bekey [28] showed that the Kalman filter equations can be written in decentralized form, allowing decomposition into smaller communicating filters at each robot. Thrun et. al [29] presented a sparse extended information filter for multi-robot SLAM, which actively removes information to ensure sparseness at the cost of approximation. With nonlinear motion and observation models, a decentralized extended Kalman filter (EKF) has an observable subspace of higher dimension than the actual nonlinear system and generates unjustified covariance reduction [30]. Huang et al. introduced observability constraints in EKF [30] and unscented smoothing [31] algorithms to ensure consistent estimation. Gao et al. [32] use random finite sets to represent landmarks at each robot and maintain a probability hypothesis density (PHD). The authors prove that geometric averaging of the robot PHDs over one-hop neighbors leads to convergence of the PHDs to a global Kullback-Leibler average, ensuring consistent maps across the robots. Zhu et al. [33] propose a distributed visual-inertial cooperative localization algorithm by leveraging covariance intersection to compensate for unknown correlations among the robots and deal with loop-closure constraints.
Compared to existing work, our contribution is to derive a fully distributed filtering approach for landmark-based SLAM from a constrained variational inference perspective. Our formulation makes a novel connection to distributed mirror descent and enables robots to achieve landmark consensus efficiently with one-hop communication only and without sharing private trajectory information.
@@ III PROBLEM STATEMENT
The robots communicate over a network represented as a connected undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with nodes $\mathcal{V}=\{1,\ldots,n\}$ corresponding to the robots and edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ specifying robot pairs that can exchange information. For example, $(i,j)\in\mathcal{E}$ indicates that robot $i$ and $j$ can exchange information. Let $A\in\mathbb{R}^{n\times n}$ be a doubly stochastic weighted adjacency matrix of $\mathcal{G}$ such that $A_{ij}>0$ if $(i,j)\in\mathcal{E}$ and $A_{ij}=0$ otherwise. Also, let $\mathcal{N}_{i}:=\{j\in\mathcal{V}|(j,i)\in\mathcal{E}\}\cup\{i\}$ denote the set of one-hop neighbors of robot $i$. This is the set of all nodes $j\in\mathcal{V}$ that have a direct connection via an edge to node $i$ and includes node $i$ itself. We consider the following problem.
@@
Given control inputs $\mathbf{u}_{i}=[\mathbf{u}_{i,0}^{\top},\ldots,\mathbf{u}_{i,T-1}^{\top}]^{\top}$ and observations $\mathbf{z}_{i}:=[\mathbf{z}_{i,0}^{\top},\ldots,\mathbf{z}_{i,T}^{\top}]^{\top}$, each robot $i$ aims to estimate the robot states $\mathbf{s}_{i}:=[\mathbf{s}_{i,0}^{\top},\ldots,\mathbf{s}_{i,T}^{\top}]^{\top}$ collaboratively by exchanging information only with one-hop neighbors $j\in\mathcal{N}_{i}$ in the communication graph $\mathcal{G}$.
@@ IV-B Distributed mirror descent
We solve the variational inference problem in (5) using the mirror descent algorithm [7]. We use mirror descent because it includes an explicit (Bregman divergence) regularization term in the objective function that can incorporate information from one-hop neighbors [35]. This allows us to formulate a distributed version of mirror descent that enforces agreement among the landmark estimates of different robots with convergence guarantees. Mirror descent is a generalization of projected gradient descent that performs projection using a generalized distance (Bregman divergence), instead of the usual Euclidean distance, to respect the geometry of the constraint set $\mathcal{Q}_{i}$. Since $\mathcal{Q}_{i}$ is a space of PDFs, a suitable choice of Bregman divergence is the KL divergence. Starting with a prior PDF $q_{i}^{(0)}(\mathbf{s}_{i})$, the mirror descent algorithm performs the following iterations:
@@
where $\delta c_{i}/\delta q_{i}(q_{i}^{(k)})$ is the Fréchet derivative of $c_{i}(q_{i})$ with respect to $q_{i}$ evaluated at $q_{i}^{(k)}$ and $\alpha_{k}>0$ is the step size.
Note that the optimizations (6) at each robot $i$ are completely decoupled and, hence, each robot would be estimating its own density over the common landmarks $\mathbf{y}$. To make the estimation process collaborative, the regularization term $\operatorname*{KL}(q_{i}||q_{i}^{(k)})$ in (6) should require that the PDF $q_{i}$ of robot $i$ is also similar to the priors $q_{j}^{(k)}$ of its neighbors $\mathcal{N}_{i}$ rather than its own prior $q_{i}^{(k)}$ alone. In our case, the PDFs $q_{j}^{(k)}(\mathbf{s}_{j})=q_{j}^{(k)}(\mathbf{x}_{j},\mathbf{y})$ are not defined over the same set of variables since each robot $j$ is estimating its own private state $\mathbf{x}_{j}$ as well. Inspired by but different from [35], to enforce consensus only on the common state $\mathbf{y}$, the KL divergence term in (6) can be decomposed as a sum of marginal and conditional terms:
@@ IV-C Linear Gaussian case
where $\otimes$ is the Kronecker product. Denoting the density and distribution at iteration $k$ as $q_{i}^{(k)}(\mathbf{s}_{i})$ and $\mathcal{N}(\boldsymbol{\mu}_{i,(k)},\Sigma_{i,(k)})$, the distributed variational inference update in (9) is computed in the following propositions.
@@
As shown in [37, Chapter 3.3], considering only two consecutive time steps in the lifted form in (11) leads to a Kalman filter. Using the result in Proposition 3, we obtain a distributed Kalman filter that incorporates the consensus averaging step in (15). In [36, Appendix C], we also derive a version of the Kalman filter that allows correlation between the motion and measurement noise.
@@ V-B Consensus averaging
which is the same as (15) except that we only consider one time step $t$ here. Then, we need to reconstruct the new joint distribution $\mathcal{N}(\bar{\boldsymbol{\mu}}_{i,t-1},\bar{\Sigma}_{i,t-1})$. Since we store the historical camera poses as $SE(3)$ matrices, Proposition 2 can not be applied directly. Following [37, Chapter 7.3.1], we define a Gaussian distribution over a historical camera pose $\underline{T}_{i,k},\ k=t-c,\cdots,t-1$ by adding a perturbation $\epsilon_{i,k}$:
@@
where $(\cdot)^{\wedge}$ defined in [37, Chapter 7.1.2] converts from $\mathbb{R}^{6}$ to a $\mathbb{R}^{4\times 4}$ twist matrix. The estimated covariance $\Sigma_{i,t-1}$ already takes account of both the poses and landmarks. For consistency of notation with Proposition 2, we denote the mean of the pose perturbation as $\boldsymbol{\mu}^{\mathbf{x}}_{i,t-1}=\mathbf{0}_{6c}$. After averaging and reconstructing the new joint distribution, $\bar{\boldsymbol{\mu}}^{\mathbf{x}}_{i,t-1}$ may be non-zero, so we need to correct the camera poses as follows:
@@ V-C State propagation
We derive a general odometry propagation step for the MSCKF algorithm, thus not necessarily requiring IMU measurements and enabling vision-only propagation. We assume an odometry algorithm (e.g., libviso2 [38]) provides relative pose measurements $\delta T_{i,t-1}$ between the frame at time $t-1$ and that at time $t$. The state of robot $i$ is propagated as:
@@
where $Ad(\cdot)$ is the adjoint of an $SE(3)$ matrix [37, Chapter 7.1.4], $\mathbf{e}_{6n}\in\mathbb{R}^{6n+3m_{t}}$ is a vector with the $6n$-th element as $1$ and the rest as 00, and $W_{i}\in\mathbb{R}^{6\times 6}$ is the odometry measurement covariance.
@@ V-D State update
The MSCKF update step follows prior work [39]. The camera pose residual is a perturbation $\underline{\epsilon}_{i,k}$ that transforms the estimated pose $T_{i,k}$ to the true pose $\underline{T}_{i,k}$, i.e. $\underline{T}_{i,k}=T_{i,k}\exp(\underline{\epsilon}_{i,k}^{\wedge})$. The landmark residual is the difference between true position $\underline{\mathbf{p}}$ and the estimated position $\mathbf{p}$, i.e., $\tilde{\mathbf{p}}=\underline{\mathbf{p}}-\mathbf{p}$, where $\mathbf{p}$ is the position mean if the landmark is in the state or the result of triangulation if it is not. When robot $i$ receives the $k$-th geometric feature observation at time $t$, denoted as $\mathbf{z}^{g}_{i,t,k}$, we linearize the observation model around the current error state $\tilde{\mathbf{s}}_{i,t,k}$ (composed of both pose and landmark residuals) and the feature position residual $\tilde{\mathbf{p}}^{g}_{i,k}$:
@@
where $\hat{\mathbf{z}}^{g}_{i,t,k}$ is the predicted observation, $H^{\mathbf{s},g}_{i,t,k}$ and $H^{\mathbf{p},g}_{i,t,k}$ are Jacobians, and $\mathbf{v}^{g}_{i,t,k}\sim\mathcal{N}(\mathbf{0},V^{g}_{i})$ is the geometric observation noise. Then, we left-multiply by the nullspace $N_{i,t,k}$ of $H^{\mathbf{p},g}_{i,t,k}$ to remove the effect of $\tilde{\mathbf{p}}^{g}_{i,k}$:
@@
Since we allow general odometry in the propagation step, potentially obtained from visual features, there may be correlation between the motion noise and the observation noise. The correlation is denoted as
$S_{i,t,k}=\mathbb{E}[\mathbf{w}_{i,t-1}\mathbf{v}_{i,t,k}^{\top}],\;\;\mathbf{w}_{i,t-1}\sim\mathcal{N}(\mathbf{0},W_{i}).$ (25)
Concatenating $\mathbf{r}^{g,0}_{i,t,k}$, $N_{i,t,k}^{\top}H^{\mathbf{s},g}_{i,t,k}$, $N_{i,t,k}^{\top}V_{i}^{g}N_{i,t,k}$, and $S_{i,t,k}N_{i,t,k}$ for all $k$ appropriately, we get an overall geometric feature residual $\mathbf{r}^{g}_{i,t}$, Jacobian $H^{g}_{i,t}$, noise covariance $V^{g}_{i,t}$ and correlation $S_{i,t}$. Similarly, we linearize the observation model for the object $\mathbf{z}^{o}_{i,t,k}$:
@@
and concatenate $\mathbf{r}^{o}_{i,t,k}$, $H^{\mathbf{s},o}_{i,t,k}$ and $V_{i}^{o}$ for all $k$ to get an overall object observation residual $\mathbf{r}^{o}_{i,t}$, Jacobian $H^{o}_{i,t}$, and noise covariance $V^{o}_{i,t}$.
Finally, by concatenating the residuals and Jacobians of both geometric and object features, we get the overall residual $\mathbf{r}_{i,t}\!=\![\mathbf{r}^{g\top}_{i,t}\ \mathbf{r}^{o\top}_{i,t}]^{\top}$ and Jacobian $H_{i,t}\!=\![H^{g\top}_{i,t}\ H^{o\top}_{i,t}]^{\top}$. As shown in  [36, Appendix C], the Kalman gain is:
$K_{i,t}$ $=(\Sigma^{+}_{i,t}H^{\top}_{i,t}+S_{i,t})(H_{i,t}\Sigma^{+}_{i,t}H^{\top}_{i,t}+V_{i,t})^{-1},$ (26)
$V_{i,t}$ $=\operatorname*{blkdiag}(V^{g}_{i,t}+H^{g}_{i,t}S_{i,t}+S^{\top}_{i,t}H^{g\top}_{i,t},V^{o}_{i,t}),$ (27)
where $S_{i,t}$ only appears for the geometric features, $\Sigma^{+}_{i,t}$ is from the prediction step, and $K_{i,t}$ can be split to $K^{\mathbf{x}}_{i,t}$ and $K^{\mathbf{y}}_{i,t}$ related to $\mathbf{x}_{i,t}$ and $\mathbf{y}_{i,t}$ respectively. The landmark mean, camera poses, and the entire covariance are updated as: